# Also at ]Alpine Quantum Technologies GmbH, Innsbruck, Austria

### v1

@@ I Introduction
$\mathcal{E}_{t}(\rho_{0})=\!\!\sum_{\alpha,\beta=0}^{d^{2}-1}\!\!\chi^{\rm err}_{\alpha\beta}(t)E_{\alpha}\rho_{\rm id}(t)E_{\beta}^{\dagger},$ (2)
@@
$E_{\alpha}\!\in\!\mathcal{B}_{\rm P}=\frac{1}{\sqrt{d}}\big{\{}\mathbb{1}_{2},\sigma_{x},\sigma_{y},\sigma_{z}\big{\}}^{\!\!\bigotimes^{N}}\!\!\!\!.$ (3)
@@
$\chi^{\rm err}(t)\in\mathsf{Pos}(\mathbb{C}^{d^{2}}),\hskip 6.45831pt\sum_{\alpha,\beta}\chi^{\rm err}_{\alpha\beta}(t)E^{\dagger}_{\beta}E_{\alpha}=\mathbb{1}_{d}.$ (4)
@@
It comes as no surprise that realistic noise sources possess more structure than the above depolarizing channel. To begin with, one- and two-qubit gates are typically driven by distinct mechanisms, and will therefore be subject to distinct error channels. Additionally, qubits idling during gate operations will invariably be subject to a noise channel that is different from that of active qubits, though both do depend on the control in general. Moreover, the error channels for the various operations including idling will be dynamical, with time dependence arising via a number of avenues. For instance, the idle channel must account for the different waiting periods, including time allocated for non-unitary operations such as reset, or (re)cooling and ion shuttling. Additionally, including the time dependence of the process matrix is a requisite if one aims at modeling noise fluctuations typically characterized by a non-zero correlation time $\tau_{\rm c}$, which leads to colored (correlated) noise. These correlations can be responsible for the non-Markovian dynamics of the QIP, which are not captured by the oversimplifying depolarizing channel. Spatiotemporal noise correlations are indeed ubiquitous in all physical hardware platforms, and break base assumptions in much of the theory work in QEC and QCVV. One way around this has been to turn correlated noise into uncorrelated Pauli noise on average through the randomized compiling approach [23]. In some contexts this average noise channel approach is inappropriate; for example in QEC where assumptions on correlations between errors matter in every shot rather than on average.
@@ I.1 Summary of the main results
In this manuscript we present a microscopically motivated reconstruction of the dynamical error map of single-qubit gates subject to colored dephasing noise, generally yielding a non-Clifford and non-Markovian map, and compare it to depolarizing and Pauli-twirled approximations. Despite its microscopic origin, our dynamical error map requires only a single quantity that is accessible in experiments: the power spectral density of the noise on the phase of the driving field. Using this reconstruction we tackle problems (i)-(iii) listed above. In order to overcome the first two limitations, and set the stage to face the third one, we go beyond a black-box QPT toolbox that makes no assumptions about the microscopic features of the device and its noise. As we show, a more detailed microscopic modeling allows us to parameterize the dynamical error map in a way so as to infer the root cause of the specific effects of the noise on single-qubit gates. For instance, we find that certain filtered noise components are the cause of a coherent miss-rotation that affects the gate even for a perfectly-calibrated driving field. Other filtered noise components fix the overall decay of the gate fidelity and also set the weights of an effective biased noise model that includes non-Markovianity.
@@
Characterizing the errors in the desired high-fidelity regime of QIPs is made challenging by the presence of SPAM errors at similar levels to gate errors. An added benefit of estimating the full dynamical error map rather than a single snapshot is that we can explore the dynamics at arbitrarily-large times. The large-time accumulated gate error contribution of the dephasing noise can be distinguished from the smaller (constant) SPAM errors. The analytic form of the reconstructed noise model can then be evaluated at shorter times to assess the errors in the high-fidelity regime of the QIP, where non-Markovian effects become more important. In particular, we show that our analysis allows us to predict the quantum processor’s performance via e.g. gate error rates in regimes where the dynamics can be proven to be non-Markovian. These error rates are then verified against other established but resource-heavier SPAM-consistent tools or benchmarking routines which only provide partial information about the channel. In our work, we apply randomized benchmarking [32, 33] to experimental trapped-ion QIPs and find a reasonable agreement with our analytical error rates. This agreement is in spite of the underlying simplifications, e.g. our analytical predictions consider dephasing noise as the sole error source and, moreover, apply a low-order truncation of a cumulant expansion to arrive at manageable time-local master equations that can be analytically solved. We present a Pauli-twirled version of this dynamical error map which, to the best of our knowledge, provides the most accurate Pauli approximation to the error map of single-qubit gates subject to time-correlated noise to date. This channel can be used in large-scale simulations of noisy Clifford-type circuits, to e.g. predict more realistic error thresholds of QEC.
@@
This manuscript is organized as follows: Sec. II constitutes the bulk of this work, where we revisit the theoretical framework to describe effects of stochastic processes in the dynamics of a single qubit. We present a new analytical solution for the dynamical quantum map of the qubit including strictly non-Markovian effects in its evolution by combining the formalism of time-local master equations with that of filter functions and noise power spectral densities. Further, we gauge the validity of our analytical methods against quasi-exact Monte Carlo simulations of the stochastic dynamics, using as a toy model the Ornstein-Uhlenbeck process to model the frequency noise of a laser. Additionally, we discuss the effects of shot noise, and how to reconstruct snapshots of the dynamical error map from data corresponding to a finite number of shots. This analytical section ends with the expressions for the average dynamical gate error with respect to ideal unitary dynamics, and a comparison of the non-Clifford and non-Markovian error map with depolarizing and Pauli-twirled channel approximations. In Sec. III, we connect our analytical formalism with experimental data obtained on a trapped-ion QIP. Here, we abandon the idealized Ornstein-Uhlenbeck dephasing and use the specific power spectral density of the frequency noise of the laser driving the ion, which is obtained by independent experimental techniques base on self-heterodyne interferometry. We close this manuscript in Sec. IV where we comment on the relevance, applicability, and extensibility of our findings to QCVV and QIP generally.
@@ II Dynamical quantum map for single-qubit gates
When the coupling of the system to the stochastic noise process is sufficiently weak, or when this process has additional properties specified below, then the dynamics can be described via a power spectral density (PSD) (see Appendix A). We are interested in particular in the filter function formalism [38, 39, 40, 41], and its more recent adaptations [42, 43, 44], in which the dynamics of the qubit are controlled by the overlap of the noise PSDs with appropriate filter functions. The estimation of PSDs is of great interest to a number of applications, and thus a wealth of methods to extract them have been developed over the years. Some of these rely on performing measurements using qubits as sensors such as in quantum noise spectroscopy [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70], which includes sophisticated adaptations [71, 56, 72, 73] that exploit the dependence of the qubit relaxation rates on the filtered noise PSD. Other methods rely on direct measurement of the noise PSDs in components, for example the local oscillators used in the driving fields [74, 75, 76, 77].
With the wealth of work already discussed elsewhere, we in turn address the question of how to estimate the dynamical error map of single-qubit gates, extending the standard long-time approach of filtered relaxation rates to the high-fidelity regime of quantum gates in which the filters act for a shorter time, and do not resolve a single frequency component of the noise PSD. Moreover, we note that the filtered relaxation rates used in quantum noise spectroscopy focus on a particular initial state while we are interested in the full dynamical quantum map. We combine the cumulant expansions in stochastic quantum dynamics [78, 79] and the filter function formalism to go beyond these limitations, capturing the noisy dynamics for high-fidelity gates and for any initial state. We derive closed analytical expressions for the dynamical error map of single-qubit gates, showing that the different snapshots depart from error channels in the Clifford group. Moreover, we find that additional filters not accounted for previously are responsible for the non-Markovianity of the dynamical map. These results are then used to address the three challenges outlined in the introduction. In particular, we present analytical expressions for the gate infidelity in Eq. (1) that go beyond previous approximations and are valid in the regime of non-Markovian dynamics. For specificity, we focus on the experimentally relevant case where the leading source of noise in the control is colored frequency or phase fluctuations leading to dephasing.
@@ II.1 Stochastic Langevin equations and the dressed-state master equation
Using a semi-classical approximation [80], the rotating frame Hamiltonian describing a Rabi flop subject to dephasing noise is given by
$\tilde{H}(t)=\frac{1}{2}\Omega\sigma_{\phi}+\frac{1}{2}\delta\tilde{\omega}(t)\sigma_{z},$ (6)
@@
Noise of the type in Eq. (8) is generally referred to as dephasing noise. While the origin of dephasing in actual hardware is usually predominantly technical, any finite linewidth oscillator used as a drive inevitably induces dephasing [79, 82, 83]. For well-controlled systems phase fluctuations are typically larger in magnitude than the noise caused by intensity fluctuations, and persist even in the absence of intensity noise [84]. As an initial modeling of this phase noise, one may take for example stochastic models of lasers above threshold, which also appear in the context of Brownian motion [85]. For instance, in the phase diffusion model [86, 87, 88, 89], $\delta\tilde{\phi}(t)$ fluctuates according to a Wiener process or, alternatively, as the result of white-noise fluctuations in the laser frequency manifested in the detuning $\delta\tilde{\omega}_{\rm d}(t)$. As discussed in Appendix B, this model is the zero correlation-time limit of an extended phase diffusion model [89, 90, 82] that uses an Ornstein-Uhlenbeck frequency noise [91, 81], accounting for non-Lorentzian PSDs [90]. These models should be understood as simplified phenomenological models that appear as limiting cases of a more detailed description of the laser [92, 79, 82, 83], and are used in our work for numerical benchmarks of the analytic derivations. These models assume, in particular, that there is no added technical noise. Moreover, they are meant to characterize the direct laser output, which will afterwards (un)intentionally be affected by other elements, changing in this way its spatiotemporal makeup and the microscopic PSD. Nevertheless, using fairly broad assumptions about general noise properties like Gaussianity or some form of stationarity, we can progress on the tomography of the dynamical error map using a generic PSD of the noise, that is in particular including technical contributions. These PSDs are the ones that can actually be accessed by independent experimental means as mentioned in the introduction. The rest of this section aims at substantiating all of the above claims starting from the effect of the stochastic dephasing on the qubit’s dynamics.
@@
When the specific SDE governing the dephasing noise is not known a priori, alternative means to generate the noise trajectories are required. Remarkably, there are situations in which the knowledge of the auto-correlation function, or the PSD, suffice for such a goal. The auto-correlation function, or covariance, of a stochastic process is defined as the statistical average of the two-time correlation functions
$C(t,t^{\prime})=\mathbb{E}[(\delta\tilde{\omega}(t)-\overline{\delta}(t))(\delta\tilde{\omega}(t^{\prime})-\overline{\delta}(t^{\prime}))],$ (10)
where $\overline{\delta}(t)=\mathbb{E}[\delta\tilde{\omega}(t)]$ is the mean value of the process that is assumed to vanish in this work. We focus on processes that are wide-sense stationary or have independent stationary increments, for which $C(t,t^{\prime})=C(t-t^{\prime})$ [94]. In this case, the PSD of the noise $S(\omega)$, that is the power of the stochastic process per unit frequency, can be defined as
$C(t)=\int_{-\infty}^{\infty}\frac{{\rm d}\omega}{2\pi}S(\omega){\rm e}^{{\rm i}\omega t}.$ (11)
@@
$\frac{{\rm d}}{{\rm d}t}{\rho}(t)=-\int_{0}^{t}\!\text{d}t^{\prime}C(t-t^{\prime})[O_{z}(t),O_{z}(t^{\prime})\rho(t)]+\rm{H.c.},$ (12)
@@
$\begin{split}\Gamma_{i}(t)=&\int_{-\infty}^{\infty}\!\!{\text{d}\omega}\,S(\omega)\,F_{\Gamma_{i}}(\omega,\Omega,t),\\ \Delta_{i}(t)=&\int_{-\infty}^{\infty}\!\!{\text{d}\omega}\,S(\omega)\,F_{\Delta_{i}}(\omega,\Omega,t),\end{split}$ (18)
@@
$F_{\,\Gamma_{1}}(\omega,\Omega,t)\approx\frac{t}{4}(S(\Omega)+S(-\Omega))=\frac{t}{2}S(\Omega),$ (21)
where we have used that classical noise PSDs are always parity even, i.e. $S(-\Omega)=S(\Omega)$ [110]. The qubit coherence decays with an exponential that depends on the noise PSD evaluated at the Rabi frequency as follows from Eq. (92) in the appendix for sufficiently long driving times $t$. This connects directly to some of the aforementioned methods of relaxation rate quantum noise spectroscopy [49, 51, 55, 54, 61, 62]. If one starts from a qubit dressed state $\ket{\pm}$, the so-called spin-locking dynamics yield the density-matrix elements $\rho_{\pm\pm}(t)=(1+{\rm e}^{-t/T_{2,{\rm eff}}})/2$, such that the relaxation of the qubit coherences can give information about the noise PSD via a renormalized dephasing time $T_{2,{\rm eff}}=2/S(\Omega)$. By sweeping the Rabi frequency $\Omega$ of the resonant drive, one can infer $S(\omega)$ in a frequency range $\omega\in[\Omega_{\rm min},\Omega_{\rm max}]$ by monitoring the coherence decay of the qubit in this long-time limit. We emphasize that the long-time limit condition is crucial for the reconstruction of the noise PSD. Spectral leakage otherwise makes PSD estimation a complicated inversion problem [58]. Active shaping of the filter function can further aid in this regard [73].
@@
$S(\omega)=\frac{c\tau_{\rm c}^{2}}{1+(\omega\tau_{\rm c})^{2}},$ (25)
@@
Figure 2: Population and coherence dynamics for resonant Rabi oscillations under Ornstein-Uhlenbeck dephasing noise. a Decay of the population $\langle\sigma_{z}(t)\rangle$ in initial states |$\ket{0}$ and |$\ket{1}$ under noisy drive, either predicted by the numerical simulations of the stochastic Langevin equations (markers) or by the analytic dressed-state master equation approach (lines). b Same for the qubit coherences $\langle\sigma_{x}(t)\rangle$ for initial states $\ket{\pm}$. Both panels use $M_{\rm MC}=10^{4}$ noise trajectories for the stochastic simulations, using the update expressions discussed in Appendix B, with $\tau_{\rm c}=5\times 10^{-4}{\rm s}$, $c=2/(10\tau_{\rm c}^{3})$, time step $\Delta t=0.05\tau_{\rm c}$ and we consider a Rabi frequency of $\Omega=40\pi/\tau_{\rm c}$. We remark that these numbers are not related to the specific trapped-ion setting, where we will use a measured noise PSD inferred from separate experiments. The common feature is that, in both cases, the correlation time is small in comparison to the timescale of interest.
For this specific stochastic process we find excellent agreement between the stochastic (numerical) evolution of Eq. (9) and our analytical approach based on filter functions according to Eqs. (92) and (96), as can be seen in Fig. 2. The time evolution in the computational basis shows the expected Rabi oscillations (Fig. 2a) damped by dephasing noise. Both models agree quantitatively on the expected exponential damping of the coherences (Fig. 2b) which is well-captured by the spin-locking renormalized dephasing time $T_{2,{\rm eff}}=2/S(\Omega)$ at long times. Remaining discrepancies can arise from a number of factors: The analytical treatment is approximate in arriving at the dressed-state master equation (12) and in truncating the Magnus expansion to solve Eq. (95). The numerical method has inherent finite sampling errors that become more apparent in the long-time limit where the coherences are very close to zero. The agreement shown here is for a set of four cardinal states. We will show in the section below that the agreement holds for any initial state via the full dynamical error map.

### v3

@@ I Introduction
$\mathcal{E}_{t}(\rho_{0})=\!\!\sum_{\alpha,\beta=0}^{d^{2}-1}\!\!\chi^{\rm err}_{\alpha\beta}(t)E_{\alpha}^{\phantom{\dagger}}\rho_{\rm id}(t)E_{\beta}^{\dagger},$ (2)
@@
$E^{\phantom{\dagger}}_{\alpha}\!\in\!\mathcal{B}_{\rm P}=\frac{1}{\sqrt{d}}\big{\{}\mathbb{1}_{2},\sigma_{x},\sigma_{y},\sigma_{z}\big{\}}^{\!\!\bigotimes^{N}}\!\!\!\!.$ (3)
@@
$\chi^{\rm err}(t)\in\mathsf{Pos}(\mathbb{C}^{d^{2}}),\hskip 6.45831pt\sum_{\alpha,\beta}\chi^{\rm err}_{\alpha\beta}(t)E^{\dagger}_{\beta}E_{\alpha}^{\phantom{\dagger}}=\mathbb{1}_{d}.$ (4)
@@
It comes as no surprise that realistic noise sources possess more structure than the above depolarizing channel. To begin with, one- and two-qubit gates are typically driven by distinct mechanisms, and will therefore be subject to distinct error channels. Additionally, qubits idling during gate operations will invariably be subject to a noise channel that is different from that of active qubits, though both do depend on the control in general. Moreover, the error channels for the various operations including idling will be dynamical, with time dependence arising via a number of avenues. For instance, the idle channel must account for the different waiting periods, including time allocated for non-unitary operations such as reset, or (re)cooling and ion shuttling. Additionally, including the time dependence of the process matrix is a requisite if one aims at modeling noise fluctuations typically characterized by a non-zero correlation time $\tau_{\rm c}$, which leads to colored (correlated) noise. These correlations can be responsible for the non-Markovian dynamics of the QIP, which are not captured by the oversimplifying depolarizing channel. Spatio-temporal noise correlations are indeed ubiquitous in all physical hardware platforms, and break base assumptions in much of the theory work in QEC and QCVV. One way around this has been to turn correlated noise into uncorrelated Pauli noise on average through the randomized compiling approach [23]. In some contexts this average noise channel approach is inappropriate; for example in QEC where assumptions on correlations between errors matter in every shot rather than on average.
@@ I.1 Summary of the main results
In this manuscript we present a microscopically motivated reconstruction of the dynamical error map of single-qubit gates subject to colored dephasing and amplitude noise, sometimes refereed to as multi-axis or universal noise in the literature. This yields a non-Clifford and non-Markovian map, which is compared to depolarizing and Pauli-twirled approximations. Despite its microscopic origin, our dynamical error map requires only a single quantity that is accessible in experiments: the power spectral of the noisy parameter of the driving field, that is phase, amplitude, or both. Using this reconstruction we tackle problems (i)-(iii) listed above. In order to overcome the first two limitations, and set the stage to face the third one, we go beyond a black-box QPT toolbox that makes no assumptions about the microscopic features of the device and its noise. As we show, a more detailed microscopic modeling allows us to parameterize the dynamical error map in a way so as to infer the root cause of the specific effects of the noise on single-qubit gates. For instance, we find that certain filtered noise components are the cause of a coherent misrotation that affects the gate even for a perfectly calibrated driving field. Other filtered noise components fix the overall decay of the gate fidelity and also set the weights of an effective biased noise model that includes non-Markovianity. In the situation where simultaneous amplitude and phase noise are not cross-correlated we find that amplitude noise manifests as simple corrections to the effective parameters describing the typically dominating phase noise.
@@
Characterizing the errors in the desired high-fidelity regime of QIPs is made challenging by the presence of SPAM errors at similar levels to gate errors. An added benefit of estimating the full dynamical error map rather than a single snapshot is that we can explore the dynamics at arbitrarily-large times. The large-time accumulated gate error contribution of the noise can be distinguished from the smaller (constant) SPAM errors. The analytic form of the reconstructed noise model can then be evaluated at shorter times to assess the errors in the high-fidelity regime of the QIP, where non-Markovian effects become more important. In particular, we show that our analysis allows us to predict the quantum processor’s performance via e.g. gate error rates in regimes where the dynamics can be proven to be non-Markovian. These error rates are then verified against other established but resource-heavier SPAM-consistent tools or benchmarking routines which only provide partial information about the channel. In our work, we apply randomized benchmarking [32, 33] to experimental trapped-ion QIPs and find a reasonable agreement with our analytical error rates. This agreement is in spite of the underlying simplifications, e.g. our analytical predictions consider dephasing and amplitude noise with no cross-correlation as the error sources and, moreover, apply a low-order truncation of a cumulant expansion to arrive at manageable time-local master equations that can be analytically solved. We present a Pauli-twirled version of the dynamical error maps, which, to the best of our knowledge, provides the most accurate Pauli approximation to the error map of single-qubit gates subject to time-correlated noise to date. This channel can be used in large-scale simulations of noisy Clifford-type circuits, to e.g. predict more realistic error thresholds of QEC.
@@
This manuscript is organized as follows: Sec. II constitutes the bulk of this work, where we revisit the theoretical framework to describe effects of stochastic processes in the dynamics of a single qubit. We present a new analytical solution for the dynamical quantum map of the qubit including strictly non-Markovian effects in its evolution by combining the formalism of time-local master equations with that of filter functions and noise power spectral densities. Further, we gauge the validity of our analytical methods against quasi-exact Monte Carlo simulations of the stochastic dynamics, using as a toy model the Ornstein-Uhlenbeck process to model the frequency noise of a laser. Additionally, we discuss the effects of shot noise, and how to reconstruct snapshots of the dynamical error map from data corresponding to a finite number of shots. This analytical section ends with the expressions for the average dynamical gate error with respect to ideal unitary dynamics, and a comparison of the non-Clifford and non-Markovian error map with depolarizing and Pauli-twirled channel approximations. In Sec. III, we connect our analytical formalism with experimental data obtained on a trapped-ion QIP. Here, we abandon the idealized Ornstein-Uhlenbeck for dephasing and amplitude fluctuations and use the specific power spectral density of the multi-axis noise of the laser driving the ion, which are obtained by independent experimental techniques based on self-heterodyne interferometry and direct measurement, respectively. We close this manuscript in Sec. IV where we comment on the relevance, applicability, and extensibility of our findings to QCVV and QIP generally.
@@ II Dynamical quantum map for single-qubit gates
When the coupling of the system to the stochastic noise process is sufficiently weak, or when this process has additional properties specified below, then the dynamics can be described via a power spectral density (PSD) (see Appendix A). We are interested in particular in the filter function formalism [38, 39, 40, 41], and its more recent adaptations [42, 43, 44], in which the dynamics of the qubit are controlled by the overlap of the noise PSDs with appropriate filter functions. The estimation of PSDs is of great interest to a number of applications, and thus a wealth of methods to extract them have been developed over the years. Some of them rely on performing measurements using qubits as sensors such as in quantum noise spectroscopy [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70], which includes sophisticated adaptations [71, 56, 72, 73] that exploit the dependence of the qubit relaxation rates on the filtered noise PSD. Other methods rely on direct measurement of the noise PSDs in components, for example the local oscillators used in the driving fields [74, 75, 76, 77].
With the wealth of work already discussed elsewhere, we in turn address the question of how to estimate the dynamical error map of single-qubit gates, extending the standard long-time approach of filtered relaxation rates to the high-fidelity regime of quantum gates in which the filters act for a shorter time, and do not resolve a single frequency component of the noise PSD. Moreover, we note that the filtered relaxation rates used in quantum noise spectroscopy focus on a particular initial state while we are interested in the full dynamical quantum map. We combine the cumulant expansions in stochastic quantum dynamics [78, 79] and the filter function formalism to go beyond these limitations, capturing the noisy dynamics for high-fidelity gates and for any initial state. We derive closed analytical expressions for the dynamical error map of single-qubit gates, showing that the different snapshots depart from error channels in the Clifford group. Moreover, we find that additional filters not accounted for previously are responsible for the non-Markovianity of the dynamical map. These results are then used to address the three challenges outlined in the introduction. In particular, we present analytical expressions for the gate infidelity in Eq. (1) that go beyond previous approximations and are valid in the regime of non-Markovian dynamics. For specificity, we start by focusing on the experimentally relevant case where the leading source of noise in the control is colored frequency or phase fluctuations leading to dephasing, as occurs for trapped-ion processors. After obtaining closed expressions for the effect of the dominant source of noise, we incorporate the correction effects caused by amplitude fluctuations on the driving field to give a universal description of the effect of noise on single-qubit gates. We emphasize, however, that the expressions of the noisy gates and their derivation do in no way require amplitude noise to be smaller in contribution than phase noise.
@@ II.1 Stochastic Langevin equations and the dressed-state master equation
We start by considering first only frequency noise on the drive. Using a semi-classical approximation [80], the rotating frame Hamiltonian describing a Rabi flop subject to dephasing noise is given by
$\tilde{H}(t)=\frac{1}{2}\Omega\sigma_{\phi}-\frac{1}{2}\delta\tilde{\omega}(t)\sigma_{z},$ (6)
@@
Noise of the type in Eq. (8) is generally referred to as dephasing noise. While the origin of dephasing in actual hardware is usually predominantly technical, any finite linewidth oscillator used as a drive inevitably induces dephasing [79, 82, 83]. For well-controlled systems phase fluctuations are typically larger in magnitude than the noise caused by intensity fluctuations, and persist even in the absence of intensity noise [84]. As an initial modeling of this phase noise, one may take for example stochastic models of lasers above threshold, which also appear in the context of Brownian motion [85]. For instance, in the phase diffusion model [86, 87, 88, 89], $\delta\tilde{\phi}(t)$ fluctuates according to a Wiener process or, alternatively, as the result of white-noise fluctuations in the laser frequency manifested in the detuning $\delta\tilde{\omega}_{\rm d}(t)$. As discussed in Appendix B, this model is the zero correlation-time limit of an extended phase diffusion model [89, 90, 82] that uses an Ornstein-Uhlenbeck frequency noise [91, 81], accounting for non-Lorentzian PSDs [90]. These models should be understood as simplified phenomenological models that appear as limiting cases of a more detailed description of the laser [92, 79, 82, 83], and are used in our work for numerical benchmarks of the analytic derivations. These models assume, in particular, that there is no added technical noise. Moreover, they are meant to characterize the direct laser output, which will afterwards (un)intentionally be affected by other elements, changing in this way its spatio-temporal makeup and the microscopic PSD. Nevertheless, using fairly broad assumptions about general noise properties like Gaussianity or some form of stationarity, we can progress on the tomography of the dynamical error map using a generic PSD of the noise, that is in particular including technical contributions. These PSDs are the ones that can actually be accessed by independent experimental means as mentioned in the introduction. The rest of this section aims at substantiating all of the above claims starting from the effect of the stochastic dephasing on the qubit’s dynamics.
@@
When the specific SDE governing the dephasing noise is not known a priori, alternative means to generate the noise trajectories are required. Remarkably, there are situations in which the knowledge of the auto-correlation function, or the PSD, suffice for such a goal. The auto-correlation function, or covariance, of a stochastic process (here: frequency fluctuations) is defined as the statistical average of the two-time correlation functions
$C_{\tilde{\omega}}(t,t^{\prime})=\mathbb{E}[(\delta\tilde{\omega}(t)-\overline{\delta}(t))(\delta\tilde{\omega}(t^{\prime})-\overline{\delta}(t^{\prime}))],$ (10)
where $\overline{\delta}(t)=\mathbb{E}[\delta\tilde{\omega}(t)]$ is the mean value of the process that is assumed to vanish in this work. We focus on processes that are wide-sense stationary or have independent stationary increments, for which $C_{\tilde{\omega}}(t,t^{\prime})=C_{\tilde{\omega}}(t-t^{\prime})$ [94]. In this case, the PSD of the noise $S_{\tilde{\omega}}(\omega)$, that is the power of the stochastic process per unit frequency, can be defined as
$C_{\tilde{\omega}}(t)=\int_{-\infty}^{\infty}\frac{{\rm d}\omega}{2\pi}S_{\tilde{\omega}}(\omega){\rm e}^{{\rm i}\omega t}.$ (11)
@@
$\frac{{\rm d}}{{\rm d}t}{\rho}(t)=-\int_{0}^{t}\!\text{d}t^{\prime}C_{\tilde{\omega}}(t-t^{\prime})[O_{z}(t),O_{z}(t^{\prime})\rho(t)]+\rm{H.c.},$ (12)
@@
$\begin{split}\Gamma_{i}(t)=&\int_{-\infty}^{\infty}\!\!{\text{d}\omega}\,S_{\tilde{\omega}}(\omega)\,F_{\Gamma_{i}}(\omega,\Omega,t),\\ \Delta_{i}(t)=&\int_{-\infty}^{\infty}\!\!{\text{d}\omega}\,S_{\tilde{\omega}}(\omega)\,F_{\Delta_{i}}(\omega,\Omega,t),\end{split}$ (18)
@@
$F_{\,\Gamma_{1}}(\omega,\Omega,t)\approx\frac{t}{4}(S_{\tilde{\omega}}(\Omega)+S_{\tilde{\omega}}(-\Omega))=\frac{t}{2}S_{\tilde{\omega}}(\Omega),$ (21)
where we have used that classical noise PSDs are always parity even, i.e. $S_{\tilde{\omega}}(-\Omega)=S_{\tilde{\omega}}(\Omega)$ [110]. The qubit coherence decays with an exponential that depends on the noise PSD evaluated at the Rabi frequency as follows from Eq. (105) in the appendix for sufficiently long driving times $t$. This connects directly to some of the aforementioned methods of relaxation rate quantum noise spectroscopy [49, 51, 55, 54, 61, 62]. If one starts from a qubit dressed state $\ket{\pm}$, the so-called spin-locking dynamics yield the density-matrix elements $\rho_{\pm\pm}(t)=(1+{\rm e}^{-t/T_{2,{\rm eff}}})/2$, such that the relaxation of the qubit coherences can give information about the noise PSD via a renormalized dephasing time $T_{2,{\rm eff}}=2/S_{\tilde{\omega}}(\Omega)$. By sweeping the Rabi frequency $\Omega$ of the resonant drive, one can infer $S_{\tilde{\omega}}(\omega)$ in a frequency range $\omega\in[\Omega_{\rm min},\Omega_{\rm max}]$ by monitoring the coherence decay of the qubit in this long-time limit. We emphasize that the long-time limit condition is crucial for the reconstruction of the noise PSD. Spectral leakage otherwise makes PSD estimation a complicated inversion problem [58]. Active shaping of the filter function can further aid in this regard [73].
@@
$S_{\tilde{\omega}}(\omega)=\frac{c\tau_{\rm c}^{2}}{1+(\omega\tau_{\rm c})^{2}},$ (25)
@@
Figure 2: Population and coherence dynamics for resonant Rabi oscillations under Ornstein-Uhlenbeck dephasing noise. a Decay of the population $\langle\sigma_{z}(t)\rangle$ in initial states |$\ket{0}$ and |$\ket{1}$ under noisy drive, either predicted by the numerical simulations of the stochastic Langevin equations (markers) or by the analytic dressed-state master equation approach (lines). b Same for the qubit coherences $\langle\sigma_{x}(t)\rangle$ for initial states $\ket{\pm}$. Both panels use $M_{\rm MC}=10^{4}$ noise trajectories for the stochastic simulations, using the update expressions discussed in Appendix B, with $\tau_{\rm c}=5\times 10^{-4}{\rm s}$, $c=2/(10\tau_{\rm c}^{3})$, time step $\Delta t=0.05\tau_{\rm c}$ and we consider a Rabi frequency of $\Omega=1/(5\tau_{\rm c})$. We remark that these numbers are not related to the specific trapped-ion setting, where we will use a measured noise PSD inferred from separate experiments. The common feature is that, in both cases, the correlation time is small in comparison to the timescale of interest.
For this specific stochastic process we find excellent agreement between the stochastic (numerical) evolution of Eq. (9) and our analytical approach based on filter functions according to Eqs. (105) and (109), as can be seen in Fig. 2. The time evolution in the computational basis shows the expected Rabi oscillations (Fig. 2a) damped by dephasing noise. Both models agree quantitatively on the expected exponential damping of the coherences (Fig. 2b) which is well-captured by the spin-locking renormalized dephasing time $T_{2,{\rm eff}}=2/S_{\tilde{\omega}}(\Omega)$ at long times. Remaining discrepancies can arise from a number of factors: The analytical treatment is approximate in arriving at the dressed-state master equation (12) and in truncating the Magnus expansion to solve Eq. (108). The numerical method has inherent finite sampling errors that become more apparent in the long-time limit where the coherences are very close to zero. The agreement shown here is for a set of four cardinal states. We will show in the section below that the agreement holds for any initial state via the full dynamical error map.